# Quantum teleportation implies symmetry-protected topological order

### v2

@@ 2.4 Physical teleportation
We are now equipped to define quantum teleportation in a general sense. Importantly, we distinguish this notion of physical teleportation—as described in Def. 3—from the slightly restricted class of quantum teleportation protocols we consider later on. We then consider a canonical example teleportation protocol involving the 1D cluster state Briegel and Raussendorf (2001); the corresponding circuit is depicted in Fig. 2.
@@
Physical teleportation refers to any unitary protocol $\mathcal{W}$ acting on a dilated Hilbert space $\mathcal{H}_{\rm dil}$ (14) that (i) fulfills the criteria for quantum state transfer (see Def. 1) and (ii) involves the combination of $M>1$ measurements, instantaneous classical communication, and outcome-dependent operations.
We stress that physical teleportation requires that the $k$ logical qubits be localized to $k$ physical qubits both before and after the protocol $\mathcal{W}$. This is in contrast to more general definitions of teleportation—which we refer to as “virtual teleportation”—in which the logical qubits may be spread out amongst many physical qubits, but can nonetheless be decoded using certain measurements thereof. This is the case for general teleportation protocols based on matrix product states Gross and Eisert (2007); Else et al. (2012); we relegate the treatment of virtual teleportation to future work Hong et al. (2023).
Importantly, the class of teleportation protocols described by Def. 3 is too broad to constrain. However, as far as we are aware, all known physical teleportation protocols—in which the final state is of the form $\lvert\Psi_{T}\rangle$ (7)—belong to a slightly restricted class of protocols. Before defining that class in Sec. 3, we briefly consider a canonical example protocol belonging to this class (depicted in Fig. 2 for $N=5$ qubits) based on the 1D cluster state Briegel and Raussendorf (2001) that teleports a single logical qubit.
Figure 2: Dilated quantum circuit for 1D cluster-state teleportation of a single logical qubit a distance $L=4$, along with Heisenberg evolution of the logical $X_{\rm L}$ operator. The $M=4$ Stinespring qubits are depicted in gray, the measurement channel in blue, and the outcome-dependent operations in orange (the operation on site $f=5$ is conditioned on the parity of the Stinespring qubits, indicated by white circles). The initial state is given by (20), with $\lvert\psi\rangle$ on the leftmost site, so that $X_{\rm L}(T)=X_{1}X_{3}X_{5}$ is a valid logical operator.
@@
$\lvert\Psi_{0}\rangle=\lvert\psi\rangle_{1}\otimes\lvert+++\dots\rangle_{\rm rest}\otimes\lvert\boldsymbol{0}\rangle_{\rm ss}\,,~{}~{}$ (20)
where |+delimited-|$\lvert$+$\rangle$| + ⟩ is the $+1$ eigenstate of $X$, $\lvert\boldsymbol{0}\rangle=\lvert 0\cdots 0\rangle$ is the default Stinespring state, and we take $N$ to be odd. We first apply to $\lvert\Psi_{0}\rangle$ (20) the (physical) unitary circuit
$\mathcal{U}=\prod_{j=1}^{N-1}\mathrm{CZ}_{j,j+1}\,,$ (21)
where the controlled-$Z$ gate $\mathrm{CZ}=\mathrm{diag}(1,1,1,-1)$ in the two-qubit computational ($Z$) basis. Applying $\mathcal{U}$ to $\lvert\Psi_{0}\rangle$ prepares the so-called cluster state Briegel and Raussendorf (2001).
We then measure $X$ on all sites $j=1,\dots,N-1$ (where $i=1$ and $f=N$ are the initial and final logical sites, respectively). This is represented by the measurement channel $\mathcal{M}$, corresponding to the blue-shaded region of Fig. 2, which is the product of the channels
$\mathcal{M}_{j}=P^{(+)}_{j}\otimes\widetilde{\mathds{1}}_{j}+P^{(-)}_{j}\otimes\widetilde{X}_{j}\,,~{}~{}$ (22)
where the projectors are in the $X_{j}$ basis, and the Stinespring registers are marked in light gray in Fig. 2.
Finally, we apply the outcome-dependent recovery operations. The parity of outcomes on “odd” sites $j=2n+1$ (with $j<N$) determines a final $Z_{N}$ rotation, while the parity of outcomes on “even” sites $j=2n+2$ (with $j<N)$ determines a final $X_{f}=X_{N}$ rotation, according to
$\mathcal{R}_{z}$ $=\mathds{1}_{f}\otimes\widetilde{P}^{(0)}_{\rm even}+X_{f}\otimes\widetilde{P}^{(1)}_{\rm even}$ (23a)
$\mathcal{R}_{x}$ $=\mathds{1}_{f}\otimes\widetilde{P}^{(0)}_{\rm odd}+Z_{f}\otimes\widetilde{P}^{(1)}_{\rm odd}\,,$ (23b)
where, e.g., $\widetilde{P}^{(n)}_{\rm even}$ projects onto the parity-$n$ configuration $\widetilde{Z}_{\rm even}=(-1)^{n}$ of the even Stinespring qubits. The channels above appear in the orange-shaded part of Fig. 2; they are named according to the logical operator they modify, and may be applied in either order. After applying these three steps, the initial logical state $\lvert\psi\rangle$ appears on site $f$ for any sequence of measurement outcomes.
Having defined the cluster-state teleportation protocol $\mathcal{W}$, we can evolve the final-state logical operator $X_{{\rm L}}=X_{f}$ backwards in time via the Heisenberg picture. We have
$\mathcal{R}^{\dagger}X_{f}\mathcal{R}$ $=\mathcal{R}^{\dagger}_{x}X_{f}\mathcal{R}_{x}=X_{f}\prod_{j=1}^{L/2}\widetilde{Z}_{2j-1}$ (24a)
$\mathcal{R}^{\dagger}Z_{f}\mathcal{R}$ $=\mathcal{R}^{\dagger}_{z}Z_{f}\mathcal{R}_{z}=Z_{f}\,\prod_{j=1}^{L/2}\widetilde{Z}_{2j}\,,~{}~{}$ (24b)
@@
$\mathcal{M}^{\dagger}\mathcal{R}^{\dagger}X_{f}\mathcal{R}\mathcal{M}$ $=X_{f}\,\prod_{j=1}^{L/2}X_{2j-1}\,\widetilde{Z}_{2j-1}$ (25a)
$\mathcal{M}^{\dagger}\mathcal{R}^{\dagger}Z_{f}\mathcal{R}\mathcal{M}$ $=Z_{f}\,\prod_{j=1}^{L/2}X_{2j}\,\widetilde{Z}_{2j}\,,~{}~{}$ (25b)
@@
$\mathcal{W}^{\dagger}X_{f}\mathcal{W}$ $=X_{i}\,\underbrace{\prod_{j=1}^{L/2}X_{2j+1}\,\widetilde{Z}_{2j-1}}_{\mathsf{S}_{x}}$ (26a)
$\mathcal{W}^{\dagger}Z_{f}\mathcal{W}$ $=Z_{i}\,\underbrace{\prod_{j=1}^{L/2}X_{2j}\,\widetilde{Z}_{2j}}_{\mathsf{S}_{z}}\,,~{}~{}$ (26b)
from which we identify $\mathsf{S}_{x}$ and $\mathsf{S}_{z}$ according to Prop. 1, and confirm that they indeed stabilize $\lvert\Psi_{0}\rangle$ (20).
@@
$\mathcal{U}\,\mathsf{S}_{x}\,\mathcal{U}^{\dagger}$ $=Z_{2}\left(\prod_{j=2}^{L/2}X_{2j-1}\right)X_{N}$ (27a)
$\mathcal{U}\,\mathsf{S}_{z}\,\mathcal{U}^{\dagger}$ $=Z_{1}X_{2}\left(\prod_{j=2}^{L/2}X_{2j}\right)Z_{N},~{}~{}$ (27b)
up to $\widetilde{Z}$ operators, which we recognize as the string order parameters of the cluster state Briegel and Raussendorf (2001). Observe that, while these operators necessarily commute in the bulk of the chain, the endpoint operators anticommute—i.e., $\{Z_{2},Z_{1}X_{2}\}=\{X_{N},Z_{N}\}=0$. This anticommutation demonstrates the nontrivial SPT order of the cluster state, as we discuss in further detail in Sec. 5.1.
@@ 2.5 Bounds on teleportation
Before defining a restricted class of teleportation protocols in Sec. 3, we briefly discuss general bounds on generic useful quantum tasks that derive from locality. We first consider bounds for unitary protocols $\mathcal{W}=\mathcal{U}$ (e.g., state transfer) that act only on $\mathcal{H}_{\rm ph}$ (1). In general, we regard $\mathcal{U}$ as a finite-depth quantum circuit (FDQC), though all of our results are compatible with $\mathcal{U}$ being generated by evolution under a local Hamiltonian, up to exponentially small operator tails. In either case, the task distance $L$ (13) obeys the Lieb-Robinson Theorem Lieb and Robinson (1972), which we state below as Theorem 2.
Theorem 2 (Lieb Robinson Lieb and Robinson (1972)).
Consider a unitary $\mathcal{W}$ realizing time evolution on $\mathcal{H}_{\rm ph}$ (1) for total time $T$, and two operators $A$ and $B$ with support in regions $X$ and $Y$, where $\lVert A\rVert=\lVert B\rVert=1$ (11) and $d(X,Y)=L$. Then,
@@
Theorem 2 has been proven in the literature for numerous classes of local Hamiltonians (and circuits) Lieb and Robinson (1972); Hastings (2010); Wang and Hazzard (2020). It has also been extended to long-range interactions Foss-Feig et al. (2015); Chen and Lucas (2019); Kuwahara and Saito (2020); Tran et al. (2020) and even local Lindblad dynamics Poulin (2010), up to minor modifications to the Lieb-Robinson velocity $v$. In the case of dynamics generated by local quantum circuits, the linear “light cone” is sharp, in that the commutator norm (28) is identically zero until $vT\geq L$, after which it is $\mathsf{O}\left(1\right)$. Detailed reviews of Lieb-Robinson bounds and their applications appear in, e.g., Refs. 48 and 54.
In the context of teleportation, the use of projective measurements and outcome-dependent operations far from the corresponding measurement can realize a task distance $L$ (13) that exceeds the Lieb-Robinson bound (28). However, strictly local nonunitary channels are captured by Lindblad dynamics, and obey the same Lieb-Robinson bound (28) Poulin (2010); hence, such channels are not useful to teleportation (compared to time evolution alone), and we do not consider them herein. Although teleportation can overcome the conventional Lieb-Robinson bound (28), it is nonetheless constrained by an alternative bound due to Ref. 36, which we now state in the specific context of teleporting $k$ qubits as Theorem 3.
Theorem 3 (Teleportation bound Friedman et al. (2022a)).
@@
While the teleportation bound (29) depends on the number of measurements $M$, the general bounds of Ref. 36 depend on the number of “measurement regions” $m$ (see Sec. 4.3). Teleporting a single qubit ($k=1$), e.g., obeys
@@
and we prove in Sec. 4 for all $k\geq 1$ that standard teleportation protocols require $M=2k$ outcomes per “region,” so the factor of two in the bound (29) is loose. In fact, it was conjectured in Ref. 36 that a minimum of two distinct measurements are required per logical qubit, per measurement region. By comparison, generating long-range entanglement—e.g., preparing an $N$-qubit Greenberger-Horne-Zeilinger (GHZ) state Greenberger et al. (1989); Piroli et al. (2021)—may succeed with a single measurement per region.
The bounds of Ref. 36 apply to generic quantum tasks that transfer or generate quantum information, entanglement, and/or correlations using arbitrary local quantum channels and instantaneous classical communication. Even when gates are conditions on the outcomes of arbitrarily distant measurements, a finite speed limit emerges. While all such protocols obey the generalized bound (30), only protocols involving local time evolution, projective measurements, and outcome-dependent feedback are compatible with violations of the standard Lieb-Robinson bound (28) Poulin (2010); Friedman et al. (2022a)—and thus, teleportation.

### v3

@@ 2.4 Physical teleportation
We are now equipped to define quantum teleportation in a general sense. Importantly, we distinguish this notion of physical teleportation—as described in Def. 3—from the slightly restricted class of quantum teleportation protocols we consider later on. We then consider a canonical example teleportation protocol involving the 1D cluster state [43]; the corresponding circuit is depicted in Fig. 2.
@@
Physical teleportation refers to any unitary protocol $\mathcal{W}$ acting on a dilated Hilbert space $\mathcal{H}^{\,}_{\rm dil}$ (14) that (i) fulfills the criteria for quantum state transfer (see Def. 1) and (ii) involves the combination of $M>1$ measurements, instantaneous classical communication, and outcome-dependent operations.
We stress that physical teleportation requires that the $k$ logical qubits be localized to $k$ physical qubits both before and after the protocol $\mathcal{W}$. This is in contrast to more general definitions of teleportation—which we refer to as “virtual teleportation”—in which the logical qubits may be spread out amongst many physical qubits, but can nonetheless be decoded using certain measurements thereof. This is the case for general teleportation protocols based on matrix product states [18, 23]; we relegate the treatment of virtual teleportation to future work [54].
Importantly, the class of teleportation protocols described by Def. 3 is too broad to constrain. However, as far as we are aware, all known physical teleportation protocols—in which the final state is of the form $\lvert\Psi^{\,}_{T}\rangle$ (7)—belong to a slightly restricted class of protocols. Before defining that class in Sec. 3, we briefly consider a canonical example protocol belonging to this class (depicted in Fig. 2 for $N=5$ qubits) based on the 1D cluster state [43] that teleports a single logical qubit.
Figure 2: Dilated quantum circuit for 1D cluster-state teleportation of a single logical qubit a distance $L=4$, along with Heisenberg evolution of the logical $X_{\rm L}$ operator. The $M=4$ Stinespring qubits are depicted in gray, the measurement channel in blue, and the outcome-dependent operations in orange (the operation on site $f=5$ is conditioned on the parity of the Stinespring qubits, indicated by white circles). The initial state is given by Eq. 20, with $\lvert\psi\rangle$ on the leftmost site, so that $X_{\rm L}(T)=X_{1}X_{3}X_{5}$ is a valid logical operator.
@@
$\lvert\Psi^{\vphantom{*}}_{0}\rangle=\lvert\psi\rangle^{\vphantom{\prime}}_{1}\otimes\lvert+++\dots\rangle^{\vphantom{\prime}}_{\rm rest}\otimes\lvert\boldsymbol{0}\rangle^{\vphantom{\prime}}_{\rm ss}\,,~{}~{}$ (20)
where |+delimited-|$\lvert$+$\rangle$| + ⟩ is the $+1$ eigenstate of $X^{\vphantom{*}}$, $\lvert\boldsymbol{0}\rangle=\lvert 0\cdots 0\rangle$ is the default Stinespring state, and we take $N$ to be odd. We first apply to $\lvert\Psi^{\,}_{0}\rangle$ (20) the (physical) unitary circuit
$\mathcal{U}=\prod_{j=1}^{N-1}\mathrm{CZ}^{\vphantom{*}}_{j,j+1}\,,$ (21)
where the controlled-$Z^{\vphantom{*}}$ gate $\mathrm{CZ}=\mathrm{diag}(1,1,1,-1)$ in the two-qubit computational ($Z^{\vphantom{*}}$) basis. Applying $\mathcal{U}$ to $\lvert\Psi^{\vphantom{*}}_{0}\rangle$ prepares the so-called cluster state [43].
We then measure $X^{\vphantom{*}}$ on all sites $j=1,\dots,N-1$ (where $i=1$ and $f=N$ are the initial and final logical sites, respectively). This is represented by the measurement channel $\mathcal{M}$, corresponding to the blue-shaded region of Fig. 2, which is the product of the channels
$\mathcal{M}^{\vphantom{\dagger}}_{j}=P^{(+)}_{j}\otimes\widetilde{\mathds{1}}^{\vphantom{*}}_{j}+P^{(-)}_{j}\otimes\widetilde{X}^{\vphantom{*}}_{j}\,,~{}~{}$ (22)
where the projectors are in the $X^{\vphantom{*}}_{j}$ basis, and the Stinespring registers are marked in light gray in Fig. 2.
Finally, we apply the outcome-dependent recovery operations. The parity of outcomes on “odd” sites $j=2n+1$ (with $j<N$) determines a final $Z^{\vphantom{*}}_{N}$ rotation, while the parity of outcomes on “even” sites $j=2n+2$ (with $j<N)$ determines a final $X^{\vphantom{*}}_{f}=X^{\vphantom{*}}_{N}$ rotation, according to
$\mathcal{R}^{\vphantom{\dagger}}_{z}$ $=\mathds{1}^{\vphantom{\dagger}}_{f}\otimes\widetilde{P}^{(0)}_{\rm even}+X^{\vphantom{*}}_{f}\otimes\widetilde{P}^{(1)}_{\rm even}$ (23a)
$\mathcal{R}^{\vphantom{\dagger}}_{x}$ $=\mathds{1}^{\vphantom{\dagger}}_{f}\otimes\widetilde{P}^{(0)}_{\rm odd}+Z^{\vphantom{*}}_{f}\otimes\widetilde{P}^{(1)}_{\rm odd}\,,$ (23b)
where, e.g., $\widetilde{P}^{(n)}_{\rm even}$ projects onto the parity-$n$ configuration $\widetilde{Z}^{\vphantom{*}}_{\rm even}=(-1)^{n}$ of the even Stinespring qubits. The channels above appear in the orange-shaded part of Fig. 2; they are named according to the logical operator they modify, and may be applied in either order. After applying these three steps, the initial logical state $\lvert\psi\rangle$ appears on site $f$ for any sequence of measurement outcomes.
Having defined the cluster-state teleportation protocol $\mathcal{W}$, we can evolve the final-state logical operator $X^{\vphantom{*}}_{{\rm L}}=X^{\vphantom{*}}_{f}$ backwards in time via the Heisenberg picture. We have
$\mathcal{R}^{\dagger}X^{\vphantom{*}}_{f}\mathcal{R}$ $=\mathcal{R}^{\dagger}_{x}X^{\vphantom{*}}_{f}\mathcal{R}^{\vphantom{\dagger}}_{x}=X^{\vphantom{*}}_{f}\prod_{j=1}^{L/2}\widetilde{Z}^{\vphantom{*}}_{2j-1}$ (24a)
$\mathcal{R}^{\dagger}Z^{\vphantom{*}}_{f}\mathcal{R}$ $=\mathcal{R}^{\dagger}_{z}Z^{\vphantom{*}}_{f}\mathcal{R}^{\vphantom{\dagger}}_{z}=Z^{\vphantom{*}}_{f}\,\prod_{j=1}^{L/2}\widetilde{Z}^{\vphantom{*}}_{2j}\,,~{}~{}$ (24b)
@@
$\mathcal{M}^{\dagger}\mathcal{R}^{\dagger}X^{\vphantom{*}}_{f}\mathcal{R}\mathcal{M}$ $=X^{\vphantom{*}}_{f}\,\prod_{j=1}^{L/2}X^{\vphantom{*}}_{2j-1}\,\widetilde{Z}^{\vphantom{*}}_{2j-1}$ (25a)
$\mathcal{M}^{\dagger}\mathcal{R}^{\dagger}Z^{\vphantom{*}}_{f}\mathcal{R}\mathcal{M}$ $=Z^{\vphantom{*}}_{f}\,\prod_{j=1}^{L/2}X^{\vphantom{*}}_{2j}\,\widetilde{Z}^{\vphantom{*}}_{2j}\,,~{}~{}$ (25b)
@@
$\mathcal{W}^{\dagger}X^{\vphantom{*}}_{f}\mathcal{W}$ $=X^{\vphantom{*}}_{i}\,\underbrace{\prod_{j=1}^{L/2}X^{\vphantom{*}}_{2j+1}\,\widetilde{Z}^{\vphantom{*}}_{2j-1}}_{\mathsf{S}_{x}}$ (26a)
$\mathcal{W}^{\dagger}Z^{\vphantom{*}}_{f}\mathcal{W}$ $=Z^{\vphantom{*}}_{i}\,\underbrace{\prod_{j=1}^{L/2}X^{\vphantom{*}}_{2j}\,\widetilde{Z}^{\vphantom{*}}_{2j}}_{\mathsf{S}_{z}}\,,~{}~{}$ (26b)
from which we identify $\mathsf{S}^{\,}_{x}$ and $\mathsf{S}^{\,}_{z}$ according to Prop. 1, and confirm that they indeed stabilize $\lvert\Psi^{\,}_{0}\rangle$ (20).
@@
$\mathcal{U}\,\mathsf{S}^{\vphantom{\prime}}_{x}\,\mathcal{U}^{\dagger}$ $=Z^{\vphantom{*}}_{2}\left(\prod_{j=2}^{L/2}X^{\vphantom{*}}_{2j-1}\right)X^{\vphantom{*}}_{N}$ (27a)
$\mathcal{U}\,\mathsf{S}^{\vphantom{\prime}}_{z}\,\mathcal{U}^{\dagger}$ $=Z^{\vphantom{*}}_{1}X^{\vphantom{*}}_{2}\left(\prod_{j=2}^{L/2}X^{\vphantom{*}}_{2j}\right)Z^{\vphantom{*}}_{N},~{}~{}$ (27b)
up to $\widetilde{Z}^{\vphantom{*}}$ operators, which we recognize as the string order parameters of the cluster state [43]. Observe that, while these operators necessarily commute in the bulk of the chain, the endpoint operators anticommute—i.e., $\{Z^{\vphantom{*}}_{2},Z^{\vphantom{*}}_{1}X^{\vphantom{*}}_{2}\}=\{X^{\vphantom{*}}_{N},Z^{\vphantom{*}}_{N}\}=0$. This anticommutation demonstrates the nontrivial SPT order of the cluster state, as we discuss in further detail in Sec. 5.1.
@@ 2.5 Bounds on teleportation
Before defining a restricted class of teleportation protocols in Sec. 3, we briefly discuss general bounds on generic useful quantum tasks that derive from locality. We first consider bounds for unitary protocols $\mathcal{W}=\mathcal{U}$ (e.g., state transfer) that act only on $\mathcal{H}^{\,}_{\rm ph}$ (1). In general, we regard $\mathcal{U}$ as a finite-depth quantum circuit (FDQC), though all of our results are compatible with $\mathcal{U}$ being generated by evolution under a local Hamiltonian, up to exponentially small operator tails. In either case, the task distance $L$ (13) obeys the Lieb-Robinson Theorem [40], which we state below as Theorem 2.
Theorem 2 (Lieb Robinson [40]).
Consider a unitary $\mathcal{W}$ realizing time evolution on $\mathcal{H}^{\,}_{\rm ph}$ (1) for total time $T$, and two operators $A$ and $B$ with support in regions $X$ and $Y$, where $\lVert A\rVert=\lVert B\rVert=1$ (11) and $d(X,Y)=L$. Then,
@@
Theorem 2 has been proven in the literature for numerous classes of local Hamiltonians (and circuits) [40, 55, 56]. It has also been extended to long-range interactions [57, 58, 59, 60] and even local Lindblad dynamics [41], up to minor modifications to the Lieb-Robinson velocity $v$. In the case of dynamics generated by local quantum circuits, the linear “light cone” is sharp, in that the commutator norm (28) is identically zero until $vT\geq L$, after which it is $\mathsf{O}\left(1\right)$. Detailed reviews of Lieb-Robinson bounds and their applications appear in, e.g., Refs. 55 and 61.
In the context of teleportation, the use of projective measurements and outcome-dependent operations far from the corresponding measurement can realize a task distance $L$ (13) that exceeds the Lieb-Robinson bound (28). However, strictly local nonunitary channels are captured by Lindblad dynamics, and obey the same Lieb-Robinson bound (28) [41]; hence, such channels are not useful to teleportation (compared to time evolution alone), and we do not consider them herein. Although teleportation can overcome the conventional Lieb-Robinson bound (28), it is nonetheless constrained by an alternative bound due to Ref. 42, which we now state in the specific context of teleporting $k$ qubits as Theorem 3.
Theorem 3 (Teleportation bound [42]).
@@
While the teleportation bound (29) depends on the number of measurements $M$, the general bounds of Ref. 42 depend on the number of “measurement regions” $m$ (see Sec. 4.3). Teleporting a single qubit ($k=1$), e.g., obeys
@@
and we prove in Sec. 4 for all $k\geq 1$ that standard teleportation protocols require $M=2k$ outcomes per “region,” so the factor of two in the bound (29) is loose. In fact, it was conjectured in Ref. 42 that a minimum of two distinct measurements are required per logical qubit, per measurement region. By comparison, generating long-range entanglement—e.g., preparing an $N$-qubit Greenberger-Horne-Zeilinger (GHZ) state [62, 63]—may succeed with a single measurement per region.
The bounds of Ref. 42 apply to generic quantum tasks that transfer or generate quantum information, entanglement, and/or correlations using arbitrary local quantum channels and instantaneous classical communication. Even when gates are conditions on the outcomes of arbitrarily distant measurements, a finite speed limit emerges. While all such protocols obey the generalized bound (30), only protocols involving local time evolution, projective measurements, and outcome-dependent feedback are compatible with violations of the standard Lieb-Robinson bound (28) [41, 42]—and thus, teleportation.